Andretti Company has a single product called a Dak. The company normally produces and sells 86,000 Daks each year at a selling price of $58 per unit. The company’s unit costs at this level of activity are given below:

Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Variable selling expenses Fixed seComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 2 Req 3 Req 4A to 4C Req 4D Req 5 AssumeComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 2 Req3 Req 4A to 4C Req 4D Req 5 AssumeComplete this question by entering your answers in the tabs below. LLLLLLLLLLL LLLLLL Req 1A Req 1B Req 2 Req 3 Req 4A to 4CReq 1A Req 1B Req 2 Req3 Req 4A to 4C Req 4D Reg 5 Due to a strike in its suppliers plant, Andretti Company is unable to purReq 1A Req 1B Req 2 Req 3 Req 4A to 4C Req 4D Req 5 An outside manufacturer has offered to produce 86,000 Daks and ship them
Answer #1

Calculation of contribution margin per Dak

Selling price

58

Less: variable costs

Direct Material

9.50

Direct Labor

11

Variable Overhead

3.50

Variable Selling Expenses

4.70

Total Variable cost

28.7

Contribution Margin per Unit

29.3

1-a Financial Advantage = Additional contribution Margin - Increased expenses

=25,800*29.3– 100,000 = $655,940

1-B yes, since benefit

2.Calculation of break even price

Direct Material

9.50

Direct Labor

11

Variable Overhead

3.5

Import Duties

1.70

Selling expenses

1.70

Total variable cost

27.4

Break even price = 27.4 + 20640/25,800 = $28.2

3.Relevant cost is the variable selling expense since manufacturing cost has already been incurred i.e. $4.70 per unit

Operating level = 86,000*25%*2/12 = 3583.33 units

4-a. Contribution margin foregone = 3583*29.3 = $104,981.9

4-b Fixed cost avoided = 774,000*70%*2/12 + 387,000*20%*2/12 = $103,200

c.Advantage of closing = 103,200 – 104,981.9 = $(1,781.9) i.e. disadvantage

d.No, should not be closed

5.Calculation of avoidable cost

Direct Material

9.5

Direct Labor

11

Variable Overhead

3.50

Avoidable Fixed manufacturing overhead

2.7

Variable selling expenses avoided

1.57

Avoidable cost per unit

28.27